x=500,n=812,95% confidence


1
Expert's answer
2021-05-07T14:19:05-0400

"\hat p=\frac{x}{n}=\frac{500}{812}=0.6158."

"95\%CI=(\hat p-1.96\sqrt{\frac{\hat p(1-\hat p)}{n}}, \hat p+1.96\sqrt{\frac{\hat p(1-\hat p)}{n}})="

"=(0.6158-1.96\sqrt{\frac{0.6158(1-0.6158)}{812}}, 0.6158+1.96\sqrt{\frac{0.6158(1-0.6158)}{812}})="

"=(0.5823,0.6492)."